# Haptic bilateral teleoperation system for free-hand dental procedures

### v1

@@ Abstract
Free-hand dental procedures are typically repetitive, time-consuming and require high precision and manual dexterity. Dental robots can play a key role in improving procedural accuracy and safety, enhancing patient comfort, and reducing operator workload. However, robotic solutions for free-hand procedures remain limited or completely lacking, and their acceptance is still low. To address this gap, we develop a haptic bilateral teleoperation system (HBTS) for free-hand dental procedures. The system includes a dedicated mechanical end-effector, compatible with standard clinical tools, and equipped with an endoscopic camera for improved visibility of the intervention site. By ensuring motion and force correspondence between the operator’s actions and the robot’s movements, monitored through visual feedback, we enhance the operator’s sensory awareness and motor accuracy. Furthermore, recognizing the need to ensure procedural safety, we limit interaction forces by scaling the motion references provided to the admittance controller based solely on measured contact forces. This ensures effective force limitation in all contact states without requiring prior knowledge of the environment. The proposed HBTS is validated in a dental scaling procedure using a dental phantom. The results show that the system improves the naturalness, safety, and accuracy of teleoperation, highlighting its potential to enhance free-hand dental procedures.
@@ I Introduction
I-A Background
Robotics is rewriting the future of healthcare, emerging as a disruptive technology capable of optimizing and revolutionizing the way some medical procedures are conducted. Technological advancements enable robots to play an increasingly prominent role in the execution of complex medical procedures, where the required scale, accuracy, and precision are often beyond human capabilities. The successful application of such technology in the medical field is feeding a wave of innovation in dentistry, where robotics is changing the game [1].
Figure 1: Manual vs teleoperated dental procedures (pictorial views). In the upper right corner, is the console of The Senhance® Surgical System [2].
@@ I-A Background
The World Health Organization estimates that oral diseases affect nearly 3.5 billion people worldwide. However, access to primary oral health services is limited due to the high costs of procedures, the unequal distribution of oral health professionals, and the lack of adequate healthcare facilities [3]. Furthermore, many procedures are conducted free-hand by uninformed operators, resulting in highly inefficient and disastrous outcomes.
@@
Traditionally, routine, prosthetic, periodontal, and some surgical dental procedures are performed free-hand. Routine procedures include the treatment of dental caries and endodontic therapy, characterized by the mechanical excavation of the damaged or infected dental structure, typically utilizing burs attached to a rotary dental handpiece, followed by the replacement of the lost structure with a synthetic filling material [4]. Prosthetic procedures, on the other hand, focus on the replacement of missing or severely damaged teeth through the use of prostheses, such as inlays, crowns, or bridges. These procedures involve the preparation of the tooth, which entails the removal of hard tissue from the damaged or healthy tooth using rotary instruments mounted on dental handpieces [5]. Periodontal procedures include scaling and root planing (SRP), which involve the removal of plaque and calculus from the surfaces of the teeth and the smoothing of root surfaces, typically employing manual hand instruments and inserts mounted on sonic or ultrasonic handpieces [6]. Finally, although computer-assisted implant surgery (s-CAIS), dynamic computer-assisted implant surgery (d-CAIS), and robotic computer-assisted implant surgery (r-CAIS) have shown a significant increase in the accuracy of dental implant placement [7], many operators still prefer the free-hand method for dental implantology [8].
@@
Each of these procedures, even the simplest one, requires accuracy, manual skills and dexterity on the part of the operator. Indeed, the latter has to cope with a highly unstructured environment, which is full of natural obstacles that occlude the view of the surgical site and reduce the space for maneuvering, requiring constant changes in the orientation of the surgical tool and the operator’s posture. The successful outcome of a procedure, as well as the comfort for the patient, therefore, depend on both human factors, i.e., the experience and skills of the human operator [9], the adequacy of the instrumentation used and the duration of the treatments [10], and environmental factors, i.e., difficult anatomical conditions and less accessibility to the surgical site [11]. In particular, the latter factors can significantly affect the operator’s comfort. In [12] a high prevalence of musculoskeletal disorders among dentists is reported. The main causes of these disorders can be traced back to an incorrect application of ergonomics, such as awkward postures, repetitive and prolonged movements, use of small or narrow tool handles, exertion of forces on hard or sharp objects, and exposure to vibrations [13]. Therefore, in the process of training new operators, much emphasis is placed on learning skills related to space utilization and positioning of surgical and non-surgical instruments, e.g., mirrors, to overcome occlusions, while maintaining healthy and comfortable posture [14]. In addition to this, to ensure the safety and effectiveness of the procedure, new operators must acquire skills related to the correct amount of pressure to apply to the site of interest, depending on the type of tool used. Consequently, the sense of touch plays a key role in achieving successful procedures. This is made evident by emerging trends in teaching, which make use of virtual environments and haptic devices (HD)s for training new operators [15, 16].
@@
In contrast to humans, robots are capable of achieving accuracy, precision, efficiency and scales beyond human capabilities. For this reason, the role of dental robots is to improve dental procedures, both in terms of patient satisfaction, i.e., broader access to care, reduced treatment time, successful outcome of the procedure, fast recovery, and in terms of operator comfort, i.e., reduced workload intensity, increased visibility of the surgical field, improved ergonomics [17], as shown in Figure 1. Haptic teleoperations are a technology that combines the robot’s skills with the flexibility, dexterity, and cognitive capabilities of humans. This technology is applied in multiple domains, such as telesurgery and space exploration, as it overcomes barriers in terms of distance and hazardous environments, and scales beyond human reach. By providing multiple types of sensory information, i.e. visual and haptic feedback, these systems enable the human operator to experience telepresence, defined as the perception of interacting with the remote environment as if it were directly encountered [18]. Furthermore, by integrating different haptic feedback and guidance modalities [19, 20], and shared control approaches [21], they contribute to reducing errors and enhance the efficiency, accuracy and safety of procedures, thus extending human capabilities, while also improving operator comfort and confidence.
I-B Related works
@@
The majority of research in robotics for dentistry is focused on oral implantology [22]. Robotic oral implant surgery offers significant benefits in terms of implant accuracy in partially [23] or totally endetulous patients, because of the control of angular deviation and axial errors, also compared to a s-CAIS system [24]. Better placement accuracy, however, can be achieved at the expense of surgical efficiency, as shown in [25], which compares surgical efficiency and implant positioning accuracy among three types of robotic systems: passive, semi-active, and active robots. Such technologies also enable flapless dental implant placement, speeding up recovery time, reducing postoperative pain and swelling, and enhancing patient comfort. [26] shows how a telerobotic-assisted drilling system can enhance the accuracy of the human operator while preserving awareness of the procedure through an effective transmission of cutting forces and vibrations.
@@
On the other hand, the state-of-the-art shows considerably less research interest in other free-hand procedures [1]. In the context of prosthetic procedures, robotics can enhance the precision of manual grinding using a turbine drill during tooth preparation [27] and can also compensate for hand tremors and the lack of tactile feedback in laser ablation [28, 29], improving accuracy [30] and efficiency [31]. In endodontics, robotics can automate and enhance precision and efficiency in root canal treatments [32, 33], outperforming current approaches also in effectively addressing biological and mechanical traits associated with biofilm recalcitrance and reinfection [34]. In periodontal procedures, robotic systems have the potential to increase accuracy, reducing the risk of errors and trauma to tissues by optimizing applied force, as with other free-hand procedures. They can also improve the ergonomics for the operator, and comfort for the patient by speeding up recovery time. However, for such procedures robotic technologies are still lacking [35], and there is no evidence of such advantages.
@@
Thus, although robotics emerges as a transformative technology in the context of dental practice, there are still many limitations to consider. First of all, the quality of research is still insufficient in terms of clinical validation and technology readiness level (TRL) [36]. This is heavily influenced by the high costs of these robotic systems and the need to ensure their compatibility with existing clinical tools [37]. Second, the most widely used category of robots is fully autonomous ones. However, the use of this type of robots requires the development of state-of-the-art algorithms and advanced capabilities in perception, recognition, and exception handling [37]. In this context, the presence of a human, who must possess digital and mechanical skills in addition to clinical expertise, is still necessary not only as a supervisor but also as an operator. Considering the absence of haptic feedback in most robotic systems, which would improve the efficiency and effectiveness of procedures and increase operator acceptance [38], as well as the need to provide access to care in areas with a shortage of professionals, haptic teleoperations represent a viable solution. These systems would enable the remote execution of procedures with the accuracy and efficiency of robots while preserving human cognitive abilities. Moreover, they would allow the acquisition of human control strategies directly exercised in the robot state space, simplifying the process of transferring the knowledge from the human to the robot [39] and enabling the effective and objective evaluation of operator skills during training [40]. Finally, research tends to prioritize oral implantology over other dental procedures, despite the high number of free-hand routine procedures performed daily and patient preferences, as they tend to accept these technologies more readily when the degree of invasiveness is lower [41]. Considering these limitations, [1] recommends to shift the focus from oral implantology procedures towards less invasive ones, which are better suited as forerunners in the field, and also to invest in educational robotics research in academia as a valuable tool to introduce dental robotics and overcome the barrier of acceptance of robotic systems among future dentists.
I-C Contributions
In our previous work, we proposed an HBTS to facilitate the execution of precision tasks and increase the safety of interaction using haptic teleoperations [42]. Motivated by the limitations or complete absence of robotic solutions for the most repetitive free-hand dental procedures, and recognizing the importance of haptic teleoperations in enhancing human capabilities while preserving dexterity, cognitive skills, and procedural awareness, in this paper we present an extended version of the system. Our contribution is an HBTS for free-hand dental procedures, including:
@@ I-C Contributions
  • a control system that ensures procedural safety, improves accuracy, and preserves manual dexterity, while providing motion correspondence between the operator’s movements and those of the end-effector, as observed in the camera’s view.
We validate the proposed HBTS on a free-hand procedure, consisting of dental scaling.
@@ II-B Haptic device and robot reference frames
  • $\mathcal{R}_{e}$ is the robot’s end-effector frame (computed from forward kinematics of joint measurements), corresponding to the tool center point (TCP);
@@ III Methodology
III-A System overview
The HBTS we propose, shown in Figure 3, is based on a position-force architecture [43], coupling the leader (LR) and follower (FR) robots. The LR is an impedance HD, sensing the pose of the human operator and rendering haptic feedback according to the remote interaction between the FR and the environment. The HD is controlled by a leader controller, featuring an eye-hand coordination controller (EHCC) and a haptic feedback controller (HFC). The former is responsible for capturing the HD’s stylus poses, filtering the human tremors, and generating the corresponding robot end-effector poses, so that the robot follows the stylus motion in coordination with the visual feedback. The latter is responsible for rendering the force feedback on the HD to let the human operator perceive the remote environment. Both controllers record data throughout the entire procedure. A detailed description of the objectives and operation of the two controllers is provided in Section III-B and Section III-D, respectively. The FR is a position-controlled robot, instructed to track the HD through an admittance controller, which provides the latter with a configurable degree of compliance on its target pose with respect to the environment. Such controller has three key features. First, it ensures the stability of the FR, even in case of loss of communication, since, in case of a lack of reference from the EHCC, the controller keeps following the last one received. Second, it avoids damage to the arm structure and manipulated environment, since it adapts the received references to ensure compliant interaction with the surroundings. Finally, it limits the interaction forces by scaling the received motion references through a free-hand force limitation strategy embedded in the controller itself. The force limitation strategy is explained in Section III-C2.
@@ III-A System overview
III-B Eye-hand coordination controller
In an HBTS, visual feedback plays a crucial role in overcoming the limitations caused by distance. In teleoperated free-hand dental procedures, using an eye-in-hand camera is the most effective solution [44]. This approach enables the operator to maintain a consistently centered view of the operative field, offering a significant advantage over eye-to-hand camera configurations, where the view may be off-center with respect to the area of interest or obstructed by physical obstacles. In this context, the human operator is naturally guided by the video stream transmitted by the camera to provide commands to the FR via the HD. However, in the eye-in-hand configuration, a rotation of the tool (unavoidable, to preserve dexterity) causes a rotation of the viewing angle. This creates a mismatch between the movements expected by the operator and those actually executed by the robot in the camera view, compromising the accuracy of the task.
The da Vinci Surgical SystemTM, which requires the insertion of multiple tools to implement an eye-to-hand strategy, provides the motion correspondence between the operator’s movements and those of the end-effector, observed in the camera view, by projecting the endoscope image above the surgeon’s hands by means of mirrored overlay optics [45]. In our eye-hand coordination controller, we provide the motion correspondence at the command level, by creating a coupling between the HD’s stylus and the robot’s TCP, and appropriately converting the motion of the former into the motion of the FR. To enhance the naturalness of the teleoperation, a rotational coupling is also established between the HD’s stylus and the FR’s TCP. This includes defining the desired relative rotation ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$, and performing manually the alignment. This is done by the human operator during what we indicate as engagement phase.
III-B1 Engagement phase
@@
III-B2 Human tremors filtering

### v2

@@ Abstract
Free-hand dental procedures are typically repetitive, time-consuming and require high precision and manual dexterity. Robots can play a key role in improving procedural accuracy and safety, enhancing patient comfort, and reducing operator workload. However, robotic solutions for free-hand procedures remain limited or completely lacking, and their acceptance is still low. To address this gap, we develop a haptic bilateral teleoperation system (HBTS) for free-hand dental procedures (FH-HBTS). The system includes a dedicated mechanical end-effector, compatible with standard clinical tools, and equipped with an endoscopic camera for improved visibility of the intervention site. By ensuring motion and force correspondence between the operator’s actions and the robot’s movements, monitored through visual feedback, we enhance the operator’s sensory awareness and motor accuracy. Furthermore, recognizing the need to ensure procedural safety, we limit interaction forces by scaling the motion references provided to the admittance controller based solely on measured contact forces. This ensures effective force limitation in all contact states without requiring prior knowledge of the environment. The proposed FH-HBTS is validated in a dental scaling procedure using a dental phantom. The results show that the system improves the naturalness, safety, and accuracy of teleoperation, highlighting its potential to enhance free-hand dental procedures.
@@ I Introduction
The World Health Organization estimates that oral diseases affect nearly 3.5 billion people worldwide. However, access to primary oral health services is limited due to the high costs of procedures, the unequal distribution of oral health professionals, and the lack of adequate healthcare facilities [1].
@@
Traditionally, routine, prosthetic, periodontal, and some surgical dental procedures are performed free-hand. Routine procedures include the treatment of dental caries and endodontic therapy, characterized by the mechanical excavation of the damaged or infected dental structure, typically utilizing burs attached to a rotary dental handpiece, followed by the replacement of the lost structure with a synthetic filling material [2]. Prosthetic procedures, on the other hand, focus on the replacement of missing or severely damaged teeth through the use of prostheses, such as inlays, crowns, or bridges. These procedures involve the preparation of the tooth, which entails the removal of hard tissue from the damaged or healthy tooth using rotary instruments mounted on dental handpieces [3]. Periodontal procedures include scaling and root planing (SRP), which involve the removal of plaque and calculus from the surfaces of the teeth and the smoothing of root surfaces, typically employing manual hand instruments and inserts mounted on sonic or ultrasonic handpieces [4]. Finally, although robotic computer-assisted implant surgery (r-CAIS) have shown a significant increase in the accuracy of dental implant placement [5], many operators still prefer the free-hand method for dental implantology [6].
Figure 1: Manual (left) vs teleoperated (right) dental procedures (pictorial views).
Each of these procedures, even the simplest ones, require accuracy, extensively trained manual skills and dexterity.
Remark 1.
The oral cavity is a highly unstructured environment, which is full of natural obstacles that occlude the view of the surgical site and reduce the space for maneuvering, requiring constant changes in the orientation of the clinical tool and the operator’s posture.
@@
Therefore, the successful outcome of a procedure, as well as the comfort for the patient, depend on both human factors, i.e., the experience and skills of the human operator [7], the adequacy of the instrumentation used and the duration of the treatments [8], and environmental factors [9], as highlighted in Remark 1. In particular, the latter factors can significantly affect the operator’s comfort. In [10] a high prevalence of musculoskeletal disorders among dentists is reported. The main causes of these disorders can be traced back to an incorrect application of ergonomics, such as awkward postures, repetitive and prolonged movements, use of small or narrow tool handles, exertion of forces on hard or sharp objects, and exposure to vibrations [11]. Hence, when training new operators, much emphasis is placed on learning skills related to space utilization and positioning of surgical and non-surgical instruments, e.g., mirrors, to overcome occlusions, while maintaining healthy and comfortable posture [12]. In addition to this, to ensure the safety and effectiveness of the procedure, new operators must acquire skills related to the correct amount of pressure to apply to the site of interest, depending on the type of tool used.
Remark 2.
The sense of touch plays a key role in achieving successful and efficient procedures.
This is also apparent from emerging trends in teaching, which make use of virtual environments and haptic devices (HD)s for training new operators [13, 14].
Robots can achieve accuracy, precision, efficiency beyond human capabilities. For this reason, the role of robots is to improve dental procedures, both in terms of patient satisfaction, i.e. broader access to care, reduced treatment time, fast recovery, and in terms of operator comfort, i.e., reduced workload intensity, increased visibility of the surgical field, improved ergonomics [15], as shown in Figure 1.
I-A Related works
Haptic teleoperations are a technology that enables the remote execution of complex tasks, employing a robotic system.
Remark 3.
Haptic teleoperations combine the robot skills with the flexibility, dexterity, and cognitive capabilities of humans, while overcoming barriers in terms of distance, hazardous environments, and scales beyond human reach.
@@ I-A Related works
By providing multiple types of sensory information, i.e. visual and haptic feedback, HBTSs enable the human operator to experience telepresence, i.e. the perception of interacting with the remote environment as if it were directly encountered [16]. Furthermore, by integrating different haptic feedback and guidance modalities [17, 18], and shared control approaches [19], they contribute to reducing errors and enhance the efficiency, accuracy and safety of procedures, while also improving operator comfort and confidence. For these reasons, such technology is applied in multiple domains, such as telesurgery, space exploration, and the handling of hazardous materials in nuclear or underwater environments. However, in the context of dentistry, the adoption of such systems is still very limited.
@@
The majority of research in robotics for dentistry is focused on oral implantology [20]. Robotic oral implant surgery offers significant benefits in terms of implant accuracy in partially [21] or totally edentulous patients, also compared to computer-assisted implant surgery (s-CAIS) systems [22]. Better placement accuracy, however, can be achieved at the expense of surgical efficiency, as shown in [23]. Such technologies also enable flapless dental implant placement, speeding up recovery time, reducing postoperative pain, and enhancing patient comfort [24]. [25] shows how a telerobotic-assisted drilling system can enhance the accuracy of the human operator while preserving awareness of the procedure through an effective transmission of cutting forces and vibrations.
@@
On the other hand, the state-of-the-art shows considerably less research effort in other free-hand procedures [26]. In the context of prosthetic procedures, robotics can enhance the precision of manual grinding using a turbine drill during tooth preparation [27] and can also compensate for hand tremors and the lack of tactile feedback in laser ablation [28, 29], improving accuracy [30] and efficiency [31]. In endodontics, robotics can automate and enhance precision and efficiency in root canal treatments [32, 33], outperforming current approaches also in effectively addressing biological and mechanical traits associated with biofilm recalcitrance and reinfection [34]. In periodontal procedures, robotic systems have the potential to increase accuracy, reducing the risk of errors and trauma to tissues by optimizing applied forces, as with other free-hand procedures. They can also improve the ergonomics for the operator, and comfort for the patient by speeding up recovery time. However, for such procedures, robotic technologies are still lacking [35], and there is no empirical evidence of such advantages.
@@
Thus, although robotics shows promise in dental practice, several limitations remain. First, the quality of research is still insufficient in terms of clinical validation and technology readiness level (TRL) [36]. This is heavily influenced by the high costs of these robotic systems and the need to ensure their compatibility with existing clinical tools [37]. Second, the most widely used robots are fully autonomous, which require the development of state-of-the-art algorithms and advanced capabilities in perception, recognition, and exception handling [37]. In this context, the presence of a human is still necessary not only as a supervisor but also as an operator. In light of Remark 2 and Remark 3, and considering the impact on the operator acceptance [38], haptic teleoperations represent a viable solution. Moreover, the HBTSs would allow the acquisition of human control strategies directly exercised in the robot state space, simplifying the process of transferring the knowledge from the human to the robot [39] and enabling the effective and objective evaluation of operator skills during training [40]. Finally, research tends to prioritize oral implantology over other dental procedures, despite the high number of daily free-hand routine procedures and the less invasiveness, entailing better patient acceptance [41]. Considering these limitations, [26] recommends to shift the focus from oral implantology procedures towards less invasive ones, which are better suited as forerunners in the field, and also to invest in educational robotics research in academia as a valuable tool to overcome the barrier of acceptance of robotic systems among future dentists.
I-B Contributions
Motivated by the limitations or complete absence of robotic solutions for the most repetitive free-hand dental procedures, and in light of Remark 3, in this paper we present a HBTS for free-hand dental procedures (FH-HBTS), including:
@@ I-B Contributions
  • a control system that ensures procedural safety, improves accuracy, and preserves manual dexterity, while providing motion correspondence between the operator’s movements and those of the end-effector, as observed in the camera view.
To the best of the authors’ knowledge, this is the only HBTS designed for free-hand dental procedures. Key contributions are:
  • a control system to ensure accurate control of the tool center point (TCP) and to guarantee motion correspondence, using a novel eye-hand coordination strategy;
  • a force limitation strategy that makes no previous assumption or estimation of the environment geometry, therefore accounting for the challenges outlined in Remark 1;
  • eye-hand force correspondence provided with saturation and filtering via a deadband function;
  • TRL-4 validation of the FH-HBTS on a free-hand procedure, consisting of dental scaling.
@@ II-B Haptic device and robot reference frames
  • $\mathcal{R}_{e}$ is the robot’s end-effector frame (computed from forward kinematics of joint measurements), corresponding to the TCP;
@@ III Methodology
The FH-HBTS we propose, shown in Figure 3, is based on a position-force architecture [42], coupling the leader (LR) and follower (FR) robots. The LR is an impedance HD, sensing the pose of the human operator and rendering haptic feedback according to the remote interaction between the FR and the environment. The HD is controlled by a leader controller, featuring an eye-hand coordination controller (EHCC) and a haptic feedback controller (HFC). The former is responsible for capturing the HD’s stylus poses, filtering the human tremors, and generating the corresponding robot end-effector poses, so that the robot follows the stylus motion in coordination with the visual feedback. The latter is responsible for rendering the force feedback on the HD to let the human operator perceive the remote environment. Both controllers record data throughout the entire procedure. The FR is a position-controlled robot, instructed to track the HD through an admittance controller, which provides the latter with a configurable degree of compliance on its target pose with respect to the environment. Such controller has three key features. First, it ensures the stability of the FR, even in case of loss of communication, since, in case of a lack of reference from the EHCC, the controller keeps following the last one received. Second, it avoids damage to the arm structure and manipulated environment, since it adapts the received references to ensure compliant interaction with the surroundings. Finally, it limits the interaction forces by scaling the received motion references through a free-hand force limitation strategy embedded in the controller itself.
With reference to Figure 3, depicting the complete system, the following sections examine its individual subcomponents in detail. In Section III-A and Section III-C, a detailed description of the objectives and operation of the EHCC and HFC controllers is provided. In Section III-B the robot interaction control is described.
@@
III-A Eye-hand coordination controller
In a HBTS, visual feedback plays a crucial role in overcoming the limitations caused by distance. In teleoperated free-hand dental procedures, using an eye-in-hand camera is the most effective solution [43]. This approach enables the operator to maintain a consistently centered view of the operative field, offering a significant advantage over eye-to-hand camera configurations, where the view may be off-center with respect to the area of interest or obstructed by physical obstacles. In this context, the human operator is naturally guided by the video stream transmitted by the camera to provide commands to the FR via the HD. However, in the eye-in-hand configuration, a rotation of the tool (unavoidable, to preserve dexterity) causes a rotation of the viewing angle. This creates a mismatch between the movements expected by the operator and those actually executed by the robot in the camera view, compromising the accuracy of the task.
The da Vinci Surgical SystemTM, which requires the insertion of multiple tools to implement an eye-to-hand strategy, thereby increasing the bulk of the equipment and the overall costs, provides the motion correspondence between the operator’s movements and those of the end-effector, observed in the camera view, by projecting the endoscope image above the surgeon’s hands by means of mirrored overlay optics [44]. In our eye-hand coordination controller, we provide the motion correspondence at the command level, by creating a coupling between the HD’s stylus and the robot’s TCP, and appropriately converting the motion of the former into the motion of the FR. To enhance the naturalness of the teleoperation, a rotational coupling is also established between the HD’s stylus and the FR’s TCP. This includes defining the desired relative rotation ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$, and performing manually the alignment. This is done by the human operator during what we indicate as the engagement phase.
The latter is presented in Section III-A1, while the human tremors filtering and the eye-hand coordination mapping, implementing the motion correspondence, are discussed in Section III-A2 and Section III-A3, respectively.
III-A1 Engagement phase
@@
III-A2 Human tremors filtering